In a particular experiment, a 12.2 g sample of N2 reacts completely. The mass of Mg3N2 produced is.
1
Expert's answer
2020-06-12T07:27:16-0400

3Mg+N2=Mg3N2


"n(N2)=\frac{m}{M}=\frac{12.2}{28.014}=0.43"


N2 in shortage


n(Mg3N2)=0.43


"m=M\times M=0.43\times100.9290=43.40"